There are twenty-seven animals in a barn. Some are chickens some are cows. There are 90 legs in all. How many of each animal are in the barn?
Found 3 solutions by solver91311, chungchengg, ikleyn:
Answer by solver91311(24713)



Assuming, in direct contravention of the Americans with Disabilities Act, that the chickens are all two-legged and that the cows are all four-legged:





Solve the 2X2 system for , the number of chickens, and the number of cows.

Then hope we don't get sued by some activist three-legged chicken, in Mithra's name we pray.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by chungchengg(1)

let ch = number of chickens
let co = number of cows
there are 27 animals so you can make an equation
ch + co = 27
ch = 27 - co equation # 1
since there are only 2 legs for chickens and 4 legs for cows and the total legs are 90 you can make another equation
2ch + 4co = 90 equation # 2........then substitute the equation #1 to equation # 2
2(27-co) + 4co =90
54 + 4co = 90
co = 9
since co for cows are 9...and the total animals are 27 ....chickens will be 18

Answer by ikleyn(52781)

.
There are twenty-seven animals in a barn. Some are chickens some are cows. There are 90 legs in all. How many of each animal are in the barn?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

See also the lesson
    - Problem on animals at a farm
in this site, where similar problem was solved.